Hi, can anyone help with this complex numbers problem?
In these complex numbers questions exact values for θ are to be used and answers are to be expressed as radicals:
1. If z=18(cos315°+i*sin315°) in trigonometric form find z in standard form.
2. If z=4√(3)(cos150°+i*sin150°) in trigonometric form find z in standard form.
Thank you for your time and consideration!

Answer by stanbon(75887)

1. If z=18(cos315°+i*sin315°) in trigonometric form find z in standard form.
------
x = 18*cos(315) = 18*cos(-45) = 18*cos(45) = 18[sqrt(2))/2 = 9sqrt(2)
----------------------
y = 18*sin(315) = 18*sin(-45) = 18[-sin(45)] = -18[sqrt(2)/2] = -9sqrt(2)
-----------------------------------------------------
z = (9sqrt(2))-9sqrt(2)i
=======================================
2. If z=4√(3)(cos150°+i*sin150°) in trigonometric form find z in standard form.
---
Go thru the same process.
Remember cos(150) = -cos(30) = -sqrt(3)/2
and sin(150) = sin(30) = 1/2
==============================
Cheers,
Stan H.